### v2

# MASS: Motion-Aware Spatial–temporal Grounding for Physics Reasoning and Comprehension in Vision-Language Models

Abstract
Vision Language Models (VLMs) perform well on standard video tasks but struggle with physics-related reasoning that involves motion dynamics and spatial interactions. We present a novel approach to address this gap by translating physical-world context cues into interpretable representations aligned with VLMs’ perception, comprehension, and reasoning. We present MASS, a model-agnostic approach that injects spatial–temporal signals into the VLM language space via depth-based 3D encoding and visual grounding, coupled with a motion tracker for object dynamics. We also contribute a comprehensive benchmark, MASS-Bench, consisting of $4,350$ real-world and AIGC videos and $8,361$ free-form video question–answering pairs focused on physics-related comprehension tasks, with detailed annotations including visual detections and grounding over sub-segments, as well as full-sequence 3D motion tracking of entities. To strengthen cross-modal alignment and reasoning, we apply reinforcement fine-tuning to MASS. Experiments and ablations show that our refined VLMs outperform comparable and baselines of larger size, and prior state-of-the-art (SoTA) models, achieving performance comparable to close-source SoTA VLMs ($2\%$ difference) like Gemini-2.5-Flash on physics reasoning and comprehension.
1 Introduction
Vision Language Models (VLMs) demonstrate strong reasoning and comprehension in standard video tasks such as captioning [27], event recognition [31] and scene understanding [65]. However, they struggle with complex visual cues that involve intertwined 3D spatial layouts [9, 60], motion patterns [11], and temporal dynamics [69, 21]. Achieving robust physical understanding requires VLMs not only to perceive visual cues, but also to internalize real-world physical principles and commonsense expectations about object behavior [15]. Such physical reasoning, central to human-level video understanding, remains challenging, as models must connect visual evidence with underlying physical dynamics and reason about whether observed events align with or violate real-world physics.
Unlike tasks such as segmentation or object grounding, where the supervision is explicit and localized, real-world physics must be inferred from indirect and often ambiguous visual evidence. For example, an apple falling and a person standing still are both influenced by gravity but differ substantially in spatial configuration, motion patterns, and temporal structure. As a result, VLMs often fail to generalize across diverse physical processes [4, 36], a challenge further magnified by the scarcity of datasets with dense spatiotemporal and motion-level annotations [16, 38]. Without explicit supervision, models tend to memorize superficial correlations rather than develop physics-grounded reasoning [14]. Effective physics comprehension requires several demanding prerequisites, spatial and temporal understanding, motion tracking, object detection, and visual grounding, yet these components are rarely annotated in sufficient detail [33, 53].
Figure 1: Physics-Centric Video Question Answering. Physics-aware video comprehension is challenging, as VLMs must capture fine-grained spatial–temporal cues and integrate them for higher-level reasoning. MASS introduces a motion-aware spatial–temporal grounding module that explicitly encodes object motions and scene dynamics into the language space. By enriching VLMs with structured spatial, temporal, and semantic signals, MASS significantly improves downstream reasoning, including motion and action understanding, physical-process inference, and abnormality detection (e.g., identifying the counterfactual upward motion of a basketball). MASS outperforms strong SoTA models such as GPT-4o and Gemini-2.5-Flash, demonstrating robust physics comprehension and reasoning across diverse tasks.
Most VLMs are trained primarily on real-world videos, where object motion usually follows physical laws. Although such data helps models capture common motion patterns, it also reinforces strong visual and language priors: models often assume that observed events are physically plausible because they match frequent patterns in the training data. This becomes more problematic with AIGC videos, where implausible trajectories, inconsistent depth, and temporally incoherent motion are common. Prior studies show that VLMs often miss or hallucinate such physics violations [5, 36]. For example, a model may infer that an apple is falling simply because it appears near a tree, relying on prior expectations rather than visual evidence.
These challenges highlight the need for going beyond raw video pixels and high-level textual supervision. Instead of training VLMs from scratch, which is infeasible and very costly in time and computational resource, one more data-efficient way is to augment VLMs with structured spatial and temporal representations that capture fine-grained object motions, interactions, and geometry. Recent advances in motion grounding [16], visual grounding [34, 37], detailed temporal modeling [64], and action analysis [17, 23] demonstrate that smaller specialized models can provide highly accurate spatial–temporal cues. However, existing approaches rarely explore how such modules can be systematically integrated into VLMs to elevate higher-level physics reasoning. Meanwhile, prior physics-oriented datasets [36, 29, 5] primarily rely on coarse annotations, limiting their capacity to induce deep, mechanism-level understanding.
Main Results: Our approach, MASS, aims to bridge the gap between the structured dynamics of the physical world and VLMs’ perception, comprehension, and reasoning abilities. We leverage expert models’ spatial and motion representations to explicitly ground entities and encode their 3D trajectories, enabling VLMs to reason about motion, interactions, and spatial constraints rather than relying on coarse priors. To evaluate these capabilities, we design a comprehensive free-form video QA benchmark covering physics phenomena with both factual and critical-thinking questions. For post-training, we explore both supervised fine-tuning (SFT) and reinforcement fine-tuning (RFT) using Group Relative Policy Optimization (GRPO) to strengthen cross-modal alignment and reasoning under complex physical contexts. The novel contributions of our work include:
  • We propose a benchmark, MASS-Bench, of $4,350$ real and AIGC videos with $8,361$ free-form QA pairs on spatial-temporal comprehension, physics-related reasoning, and abnormality detection. It includes entity-level motion grounded annotation with visual grounding over spatial and temporal dimension and dense spatial-motion representations across videos.
  • We introduce MASS, a model-agnostic, motion-aware spatial–temporal grounding algorithm for physics reasoning and comprehension. MASS explicitly grounds and represents spatial–motion information of entities in video, integrating visual grounding with structured spatial and temporal representations through a spatiotemporal awareness module. This enables VLMs to capture and encode object dynamics that are otherwise inaccessible from raw prompt inputs.
  • We post-train VLMs using GRPO to improve their comprehension and reasoning and cross-modality alignment over physics-related video phenomena. Experiments and ablations show that our refined VLMs outperform comparable and larger baselines, and prior state-of-the-art models, by $8.7\%$ and $6.0\%$, achieving performance close to closed-source SoTA VLMs like Gemini-2.5-Flash and validating the effectiveness of our approach.
2 Related Work
Physical Reasoning and Abnormality Detection in Videos: Despite rapid progress in AIGC video generation, generated videos still often contain prompt mismatches and violations of physics or commonsense, from earlier models such as LaVIE [54], SORA [7], and CogVideoX [63] to more recent systems including VEO3 [51], Wan2.2 [52], and COSMOS [1]. This has motivated growing work on abnormality evaluation and mitigation. VideoScore [22] trains synthetic evaluators aligned with human judgment, while recent benchmarks [29, 45, 44, 15, 36, 66] support systematic evaluation of physics and commonsense reasoning. WorldScore [18] provides a unified protocol over controllability, quality, and dynamics, and CRAVE [50] studies content-rich assessment through text–temporal fusion and motion-fidelity modeling. On the mitigation side, prior work has explored physics-based cues [24], chain-of-thought reasoning in PhyT2V [59], action-centric evaluation and correction in VideoPhy-2 [5], and motion-aware plausibility judgment in TRAVL [43]. Still, reliably identifying and addressing physical and commonsense failures remains an open challenge.
Spatial, Temporal, and Motion Understanding in VLMs: Recent work has improved video understanding in VLMs along spatial, temporal, and motion dimensions. For spatial reasoning, SpatialRGPT [12], VG-LLM [68], and LayoutVLM [49] strengthen 3D perception through depth cues, video-grounded geometry, and layout optimization. For temporal modeling, SlowFast-LLaVA [58] captures both local detail and long-range dynamics, TVS [19] improves alignment with temporally relevant segments, and T* [64] studies long-video retrieval via spatial search. Other efforts focus on motion and action understanding, including OpenMixer [6], MotionSight [17], and GroundMoRe [16]. However, physics-centric reasoning remains underexplored, as it requires jointly integrating spatial, temporal, and motion cues with physical commonsense.
Video Understanding in VLMs: Recent VLMs, including Qwen-VL [3] and InternVL [70], have improved general video understanding, yet remain limited on complex spatial, temporal, and long-video reasoning [35]. LongVLM [56] improves long-video modeling through hierarchical token merging, while VideoMind [42] uses a role-based agent with Chain-of-LoRA for efficient temporal reasoning. Video-Holmes [13] highlights persistent failures on multi-segment reasoning, and EgoLife [61] supports long-context retrieval for egocentric video assistants. Other work addresses hallucination and reasoning, including MASH-VLM [2], MVoT [28], and large-scale 3D spatial VQA data in [10]. Reinforcement-based training has also shown promise, as in Video-R1 [20] and VideoChat-R1 [32].
3 MASS-Bench: A Motion-Grounded Physics Reasoning and Comprehension Benchmark
In this section, we introduce the details of our benchmark. High-quality training data is essential for enabling video models to understand the physical world, yet existing resources fall short when tasks extend beyond scene-level captioning to entity-level spatiotemporal reasoning and factual comprehension, such as physics understanding, commonsense reasoning, or abnormality detection in AIGC videos. This gap largely stems from the lack of datasets that (1) provide fine-grained, spatially and temporally consistent annotations of entities across frames, and (2) include a balanced mixture of positive examples that follow real-world physical dynamics and negative examples that intentionally violate physical laws. In the absence of such enriched and balanced supervision, VLMs remain limited to surface-level perception; by contrast, datasets meeting these criteria enable more robust, physics-aware reasoning over dynamic video content.
Figure 2: Data Exhibition of MASS-Bench. MASS-Bench provides two question types, including factual and critical-thinking, to evaluate physics-driven video understanding. For each video–question–answer pair, we supply rich motion-grounding annotations, including temporal segmentation, entity-level visual grounding, temporal profiles across the full video, and motion attributes such as first/last positions and 3D displacement vectors. These structured spatial–temporal cues transform complex physics-related perception into interpretable representations that support more reliable physical reasoning. Additional dataset details are provided in the supplementary material.
Data Collection: MASS-Bench is a free-form video question-answering dataset designed to strengthen VLMs’ physics understanding, built from real-world videos sourced from Motion-Sight [17] and ActivityNet [8], alongside AIGC videos collected from VideoPhy2 [5], VideoHallu [36], and generative models such as COSMOS [1], Wan2.2 [52], and Sora [7]. MASS-Bench consists of $4,350$ videos, composed by $8,361$ free-form video question-answering pairs related to physics-related comprehension tasks. The dataset includes both positive and negative demonstrations with detailed human annotations providing balanced coverage of labels and reasoning steps. This design enables VLMs to learn correct physical processes from real-world videos, while also recognizing artifacts and violations that commonly arise in AIGC videos but are rare in natural scenes.
Questions: As shown in Figure 2, our dataset contains two types of reasoning questions: Factual questions that assess models’ grounding of real versus hypothetical scenarios, asking whether a physical process follows real-world principles or identifying the underlying mechanism, and, for AIGC videos, determining whether the physics aligns with reality or exhibits artifact-driven abnormalities. and Critical-thinking questions that require inferring causes, intentions, or detecting physical or commonsense abnormalities beyond explicit annotations. We re-annotate multiple-choice and labeled physics tasks into open-ended QA using Claude-4-Sonnet, and all ground-truth answers are further verified by humans to ensure physical consistency and accessibility to both humans and VLMs.
Categorization: To capture the spectrum of physics-related cognition required in MASS-Bench, we group all questions into five categories ordered from basic to advanced: Spatial Understanding (SU, 33.3%), Temporal Understanding (TU, 19.5%), Motion & Action Recognition (MAR, 14.4%), Physics Comprehension (PC, 15.6%), and Physical Abnormality Detection (PA, 17.1%). VLMs must first establish spatial and temporal awareness, then recognize motion patterns, before progressing to higher-level physics comprehension and real-world violation detection. The full category definitions and distribution are provided in Table 5 of the supplementary material.
Motion Grounding. Beyond the question, video, and ground-truth answer, each pair includes motion-grounding annotations: temporal segmentation indexed by frame ranges, entity-level visual grounding with bounding boxes, temporal profiles tracking each entity across the full video, and per-segment motion attributes (first/last 3D positions and motion vectors). These cues convert physics-intensive perceptual challenges into structured textual representations for more reliable reasoning. Full annotation details are provided in the supplementary material.
Figure 3: Overview of MASS: We use a model-agnostic framework to enhance video understanding with explicit spatial and motion awareness. Beyond standard visual transformer encoders that process video inputs (e.g., LLaVA-OneVision [26], Qwen2.5-VL [3]), we introduce a visual grounding module to better align queried entities with corresponding visual cues. Depth estimation captures spatial geometry, while motion tracking captures temporal dynamics across frames. These cues are fused into entity-centric motion traces and converted into motion-derived textual tokens that are appended to the textual input stream. During post-training, we keep the spatial–temporal encoders frozen, i.e., used off-the-shelf without parameter updates, and apply reinforcement fine-tuning (RFT) only to improve the LLM backbone’s comprehension of the additional multimodal information.
4 MASS: Model-Agnostic Approach
High-quality human annotations are essential for strengthening VLMs’ understanding of physics-related visual contexts, yet data alone cannot address their core limitations in modeling spatial layouts and motion dynamics. Effective physical reasoning requires isolating relevant cues from noisy video content. While one solution is to retrain models with heavy preprocessing pipelines, this incurs substantial cost. Instead, we propose a more data-efficient alternative: augmenting VLMs with dedicated spatial and motion encoders that explicitly extract and represent key visual signals. As shown in Figure 3, our model-agnostic design combines lightweight architectural refinement with targeted training to enhance spatiotemporal perception, cross-modal alignment, and physics-aware reasoning, ultimately reducing hallucinations and improving physical comprehension.
4.1 Entity-Centric Visual Grounding
Understanding physics-related dynamics in videos often depends on accurately identifying and tracking the relevant entities across time. This remains challenging because videos contain dense spatial–temporal content, while language queries are often underspecified about which objects are relevant. Grounding is therefore critical for localizing queried entities, preserving their motion and spatial context, and filtering out irrelevant information. It also helps mitigate the information loss introduced by LLM-based VLMs such as LLaVA [40], which compress rich visual signals into a limited set of language-aligned tokens.
We begin by semantically extracting the key components from each probing question to narrow the target for visual grounding. Using Grounding-DINO [41], we detect bounding boxes for the queried entities and apply a dynamic temporal-resolution scheme that automatically selects an appropriate sub-sequence length based on the input video’s duration. This segmentation strategy increases temporal coherence within each chunk, improving object tracking and downstream processing while balancing accuracy and efficiency. It also preserves grounding consistency over time and reveals temporal artifacts such as sudden appearance or disappearance, which are common in AIGC videos. For each detected entity, we construct a time-aligned grounding profile. Finally, we apply SAM2 [46, 47] to generate segmentation masks for the grounded regions, enabling subsequent spatial–temporal analysis.
4.2 Spatial Motion Feature Extraction
Physical reasoning depends on accurate knowledge of where entities are, how they move, and how their interactions evolve over time. After identifying the key entities referenced by the probing questions, the main challenge lies in the VLM’s limited ability to comprehend their spatial positions, appearances, and dynamics across frames. Without explicit spatial–temporal understanding, VLMs struggle to track entities, recognize actions, and maintain frame-level consistency, ultimately hindering higher-level physical reasoning. To address these limitations, we extract explicit spatial–temporal representations for each queried entity by integrating domain-specialized visual encoders capable of translating raw spatial–motion cues into structured features accessible to the VLM.
Leveraging the narrowed-down, visually grounded entities identified by the probing questions, we perform motion tracking for each entity within its video sub-sequences using CoTracker3 [25], which tracks arbitrary point trajectories inside the grounded regions. For each entity, we compute the spatial deviations across time, including averaged motion vectors and the first/last tracked positions. In parallel, we estimate per-frame depth using Depth Anything V2 [62] to provide spatial awareness. By aligning sampled RGB frames with their corresponding depth maps, we project the tracked motion vectors over each tracking points throughout the entity into 3D space, yielding both 3D positions and full spatio-temporal motion trajectories for each entity.
4.3 Visual Feature Representation
After encoding visual features across different modules, the key challenge lies in integrating heterogeneous spatial–motion outputs into the VLM pipeline without disrupting its native architecture. Prior work [57, 36] shows that LLM backbones better comprehend abstract, factual, and physics-related concepts when such information is expressed in the language space rather than as raw visual features.
Motivated by this, we fuse visual cues through natural-language representations aligned with the tokenized questions. We compute these features because VLMs cannot reliably infer fine-grained spatial layouts or motion dynamics from compressed visual embeddings; providing explicit, structured cues supplies the physical information needed for stable and interpretable reasoning. Our visual-feature serialization converts spatial–motion signals into structured motion-trajectory sequences for each entity, including 3D start/end positions, motion vectors, bounding boxes within each video sub-segment, and associated temporal indices. These attributes are expressed through predefined natural-language templates (detailed in the supplementary material), enabling the VLM to access precise spatial–temporal information in the modality it handles most effectively.
4.4 Training Pipeline
Although visual grounding and spatial–motion features provide essential cues, VLMs cannot directly leverage these signals for higher-level reasoning because their representations are misaligned with those learned during pre-training. As a result, even precise motion trajectories, spatial positions, and temporal indices such as knowing an object is intact at position A at time 1 and broken at position B at time 2 cannot be meaningfully synthesized without additional alignment. The model must bridge this representational gap to integrate structured cues, filter irrelevant details, and reason over scene evolution rather than merely retrieving or describing visual content.
We perform post-training to enhance the VLM’s reasoning capabilities to bridge this gap. Chain-of-thought (CoT) reasoning [55] helps the model process complex temporal and physical concepts more effectively. By encouraging intermediate reasoning steps, CoT aligns the VLM’s decision process with the multi-stage deductions needed for physical comprehension. This training enables the VLM to operate on grounded spatial–motion features, integrate temporal changes, and produce coherent, physics-aware explanations rather than treating cues as isolated descriptors. We evaluate both Supervised Fine-tuning (SFT) and Reinforcement Fine-tuning (RFT). For RFT, we adopt Temporal Group Relative Policy Optimization (T-GRPO) [20], which adds temporal-aware rewards on top of vanilla GRPO [48], a key component for video understanding. We further incorporate ROUGE [39] for semantic coherence, a temporal reward to strengthen temporal reasoning, and a format reward to encourage proper CoT generation.
5 Empirical Results
For our experiments, we evaluate 12 widely used state-of-the-art VLMs as baselines (e.g., GPT-4o and Gemini-2.5-Flash), covering 7B–38B parameter scales. We report performance comparisons between these baselines and VLMs augmented with MASS in Table 1, and present ablations on our backbone models and post-training strategies in Table 3. For evaluation, we use GPT-4.1-Mini as an automated judge to assess the factual correctness between each model’s answer and the ground-truth annotations regarding these physical comprehension and reasoning questions provided. We provide additional ablation studies are provided in the supplementary material, including real-world video QA generalization (Section E), noise-induced robustness analysis (Section F), and inference efficiency measurements (Section G).
Model SU TU MAR PC PA Overall
Baselines
VideoLLaVA 36.39 31.76 21.01 38.55 17.76 30.72
InternVL3.5-8B-Flash 47.22 45.25 39.40 58.65 32.76 44.68
InternVL3-8B 48.20 49.89 39.29 57.20 32.47 45.59
InternVL3.5-14B-Flash 47.68 49.23 41.96 57.52 31.03 45.71
LLaVA-OneVision-7B 49.18 51.99 38.12 56.02 30.46 45.79
Qwen2.5-VL-7B 55.87 53.03 54.89 52.36 30.21 52.30
Qwen3-VL-8B 57.29 50.77 51.92 58.65 37.07 53.23
R1-finetuned Reasoning VLMs
VideoChat-R1 51.20 48.66 47.34 52.30 31.09 47.05
Video-R1 57.49 55.19 47.65 63.53 33.91 53.08
Larger / Close-Source
InternVL3.5-38B-Flash 52.19 49.88 43.29 61.45 32.71 48.71
GPT-4o 54.47 55.13 55.62 60.15 31.32 52.81
Qwen3-VL-32B 58.87 50.59 48.05 59.84 38.63 53.44
Gemini-2.5-Flash 65.82 59.65 68.57 65.15 43.80 63.18
VLMs with MASS
LLaVA-OneVision-7B + MASS 61.26 65.41 58.15 61.45 42.37 59.24
Qwen2.5-VL-7B + MASS 63.52 63.06 64.07 65.06 45.79 61.93
Table 1: Accuracy (%) across categories in MASS-Bench: Performance of VLMs across five physics-related categories—Spatial Understanding (SU), Temporal Understanding (TU), Motion & Action Recognition (MAR), Physics Comprehension (PC), and Physical Abnormality Detection (PA). Models are grouped by family and sorted by overall accuracy. Scaling to 32B–38B parameters yields only marginal gains over 8B counterparts, whereas integrating MASS with 7B models substantially improves performance and achieves accuracy comparable to Gemini-2.5-Flash by leveraging explicit spatial–temporal representations.
5.1 Baseline Insights: Why Motion-Aware Spatial-temporal Grounding Matters
Spatial–Temporal Signals Are Critical for Physics Understanding. We evaluate baseline VLMs across a wide parameter range, from 7B–8B models in the LLaVA [40], Qwen-VL [3], and InternVL [70] families, up to 32B–38B variants and close-source models such as GPT-4o and Gemini-2.5-Flash. Under the same evaluation setting, neither newer variants nor larger models consistently perform better. For instance, Qwen2.5-VL-7B and Qwen3-VL-8B achieve similar results, and InternVL3.5-8B-Flash shows no clear advantage over InternVL3-8B. This trend also appears at larger scales: InternVL3.5-38B-Flash improves by only $4.0\%$ over its 8B version, while Qwen3-VL-32B remains comparable to Qwen3-VL-8B. By contrast, incorporating MASS yields substantial improvements. Qwen2.5-VL-7B+MASS and LLaVA-OneVision-7B+MASS exceed the previous best model, Qwen3-VL-8B, by $8.7\%$ and $6.0\%$, respectively. Qwen2.5-VL-7B+MASS also matches Gemini-2.5-Flash overall and outperforms it on physics abnormality detection, the most reasoning-heavy category. Together, these results show that explicit spatial–temporal representations supply physical cues that standard VLMs often miss, and are therefore key to reliable physics reasoning.
Training Data Quality is Crucial for Reasoning. We also include two R1-finetuned VLMs, Video-R1 and VideoChat-R1, in Table 1, both trained with GRPO to enhance reasoning. However, neither model shows clear improvement on physics comprehension tasks compared with non–GRPO fine-tuned VLMs. In contrast, integrating MASS with Qwen2.5-VL-7B and LLaVA-OneVision-7B yields substantially higher performance, highlighting the importance of high-quality training data. Detailed annotations, along with both positive (successful) and negative (failure) demonstrations, are essential for strengthening reasoning and counteracting language priors when tackling out-of-domain physics tasks. Without such data, RFT alone offers limited gains and may even reinforce hallucinations arising from misinterpreting visual evidence.
5.2 Ablation Study: The Role of Reasoning
Reasoning Dominates Beyond the Base Model. Table 3 compares post-training results on two backbone VLMs, Qwen2.5-VL-7B [3] and LLaVA-OneVision-7B [26]. Across both models, enhanced reasoning contributes far more than the backbone choice itself: pairing Qwen2.5-VL-7B with MASS yields a $9.6\%$ improvement, while LLaVA-OneVision-7B combined with MASS achieves a larger $13.5\%$ gain. After post-training, the base model’s influence becomes secondary, as physics comprehension, especially challenging tasks like physics abnormality detection, demands integrated reasoning over visual physics cues, spatial–temporal structure, and commonsense priors. This synergy, rather than raw model capacity, ultimately drives the performance improvements.
Visual Evidence and Motion Context Are Complementary. We ablate the input modality using three settings: (i) Text-only (motion profile + question), (ii) Visual-only (video tokens + question), and (iii) Both (video tokens + motion profile + question, our default). Table 2 reports these results alongside the SFT vs. GRPO comparison. Text-only consistently underperforms visual-only across all models (e.g., GPT-5.2: 42.09% vs. 56.17%; Gemini-2.5-Pro: 43.39% vs. 63.41%), confirming that many questions are inherently video-dependent. While the motion profile provides useful temporal summaries, it cannot replace the spatial detail and grounded evidence in the original video. Consistent with this, about 26% of text-only outputs are judged as uncertain by the LLM judge. Using both modalities yields the best results. Under the Both setting, GRPO + MASS improves Qwen2.5-VL-7B by 9.6% and LLaVA-OneVision-7B by 13.0%. This shows that the motion profile is most effective as complementary context rather than a substitute for video. The benefit also depends on model capacity: large models such as GPT-5.2 and Gemini-2.5-Pro can exploit motion-profile text even without post-training, whereas 7B models gain little from the added text unless they are further trained to integrate it. This pattern suggests that the improvement does not come from extra textual context alone, but from learning to align motion-profile cues with visual evidence for better physics reasoning.
Why RFT Matters: SFT Alone Is Not Enough. Table 3 compares SFT and RFT strategies. SFT yields limited benefit overall: it slightly reduces performance for Qwen2.5-VL-7B by 0.3% and brings only a modest 4.2% gain for LLaVA-OneVision-7B, leaving both close to the off-the-shelf models. By contrast, GRPO produces large and consistent gains across both backbones, reaching 61.93% on Qwen2.5-VL-7B and 59.24% on LLaVA-OneVision-7B. We believe this is because GRPO optimizes task-level rewards directly, making it better suited than SFT for learning physical reasoning that depends on temporal evidence and causal consistency. The gains are especially clear on SU, TU, MAR, and PC, where stronger reasoning is needed to connect motion cues with physical dynamics. PA is the main exception, where SFT performs better in our reruns, likely because PA depends more on answer calibration: the model must not only detect the abnormality, but also describe it in a precise free-form way that matches the reference. Even so, GRPO remains the stronger strategy overall.
Input Variants Training Strategy
Model Text Visual Both SFT GRPO
GPT-5.2 42.09 56.17 59.46
Gemini-2.5-Pro 43.39 63.41 64.30
Qwen2.5-VL-7B 32.14 52.30 52.38 50.40 52.89
+ GRPO + MASS 39.24 52.94 61.93 52.01 61.93
LLaVA-OneVision-7B 31.69 45.79 46.21 47.60 50.33
+ GRPO + MASS 35.63 46.72 59.24 50.03 59.24
Table 2: Input modality ablation (left) and training strategy comparison (right) for models with and without MASS. Results are reported as accuracy (%) on MASS-Bench.
Model SU TU MAR PC PA Overall
Qwen2.5-VL-7B 55.87 53.03 54.89 52.36 30.21 52.30
+ SFT 53.68 54.32 48.10 59.02 47.54 52.01
+ GRPO 63.52 63.06 64.07 65.06 45.79 61.93
LLaVA-OneVision-7B 49.18 51.99 38.12 56.02 30.46 45.79
+ SFT 47.81 51.22 49.84 50.00 49.26 50.03
+ GRPO 61.26 65.41 58.15 61.45 42.37 59.24
Table 3: Ablation of post-training strategies. Accuracy (%) of base VLMs, SFT variants, and GRPO-enhanced models across five physics-related categories. GRPO provides substantial gains over both base models and SFT in most categories.
Training Data and Module Ablations. We ablate the GRPO training data and the motion representation under the same training budget. Table 4 shows two main takeaways. First, the balanced Both setting works best on both backbones. Real-only training brings almost no gain over the off-the-shelf models, while Syn-only helps more but still remains below Both. This suggests that the two data sources help in different ways: real videos add natural variation in appearance and motion, while synthetic data more clearly expose the physical patterns the model needs to learn. Second, 3D motion matters. Replacing the full 3D motion representation with 2D motion lowers accuracy for both models. Because this ablation only removes the depth-aware projection, the drop is not just about having fewer useful tokens; it shows that depth-aware motion provides information that 2D motion alone misses. Overall, the gains come from combining complementary training data with motion cues that better reflect the physical structure of the scene.
Baselines (MASS) Training Data Modules
Model (+ MASS) w/ w/o Real-only Syn-only 2D Motion
Qwen2.5-VL-7B 61.93 52.30 52.40 56.30 56.50
LLaVA-OneVision-7B 59.24 45.79 49.92 53.41 52.30
Table 4: Ablation of GRPO training data and depth module. Accuracy (%) on MASS-Bench. w/ and w/o denote the full MASS pipeline (real+syn data, 3D motion) and the base model without MASS, respectively. Real-only and Syn-only vary the training data, while 2D Motion removes the depth-aware 3D projection.
5.3 Discussion
The Impact of Motion-aware Spatiotemporal Grounding. In Table 1 and Table 3, we observe that across model architectures, scales, and post-training strategies, VLMs equipped with our motion-aware spatiotemporal grounding outperform their baselines and reach performance comparable to Gemini-2.5-Flash. These gains stem from explicitly integrating spatiotemporal signals, as standard VLM encoders struggle to utilize or correlate spatial and temporal cues effectively. Providing structured representations of key video dynamics enables stronger physics comprehension and reasoning while filtering irrelevant or distracting visual information.
Reasoning is Essential for Understanding Physics. Motion-aware spatiotemporal grounding is necessary but not sufficient for physics comprehension. Reasoning is the key mechanism that integrates and interprets the grounded information. As shown in Table 3, comparing SFT and GRPO reveals that even with motion-aware spatiotemporal signals, learning purely from input–output correlations (as in SFT) does not help—and often degrades—performance on physics reasoning tasks. Given the complexity and interdependence of physical processes, VLMs cannot grasp the underlying dynamics through grounding alone and may lose focus during question answering. In contrast, reasoning via chain-of-thought encourages VLMs to reflect on grounded cues and combine them with commonsense knowledge, effectively bridging the gap between visual perception and real-world physical understanding.
Challenges in Physical Abnormality Detection Remain. A key observation from Table 1 and Table 3 is that Qwen2.5-VL-7B and LLaVA-OneVision-7B integrated with MASS outperform all other VLMs, including Gemini-2.5-Flash, on physical abnormality detection. However, compared with more grounded physics comprehension tasks, abnormality detection is inherently more difficult. It requires VLMs to overcome hallucinations rooted in their learned priors about real-world physics [36] and to reason about why a visual process violates physical laws. While our method yields a notable $12\%$ improvement on this task, indicating the promise of motion-aware spatiotemporal grounding for enhancing perception and reasoning, further advances are needed to robustly handle these challenging cases.
6 Conclusion, Limitations and Future Work
We present a model-agnostic approach that bridges the gap between raw video content and physics-aware reasoning by translating physical-world cues into structured representations compatible with VLM inference. We introduce MASS-Bench, to our knowledge the first benchmark pairing positive and negative video demonstrations with QA on spatial-temporal comprehension, physics-based reasoning, and abnormality detection, supported by fine-grained visual detections, sub-segment grounding, and 3D motion tracking. Our method, MASS, injects spatial–temporal signals into the VLM language space via depth-based 3D encoding, visual grounding, and motion-aware trajectory modeling, yielding consistent gains over strong baselines and recent SoTA VLMs on both real-world and AIGC videos.
Limitations. Although post-training improves reasoning, it does not fully solve physical understanding in real-world settings. Performance may still decline in crowded scenes, where multi-object tracking becomes more demanding, and hallucinations persist in reasoning-heavy cases when subtle visual evidence is misread or overshadowed by language priors. MASS-Bench is entity-centric, focusing on the top-$K$ targets, and thus does not evaluate highly crowded scenes. Robust handling of strong camera ego-motion also remains open. Encouragingly, MASS remains reasonably stable under noisy motion cues, with gradual rather than catastrophic degradation (see Section F in Supplementary Material).
Future Work. Future work will improve long-range motion grounding and temporal reasoning in longer videos, and make multi-object tracking more scalable in crowded scenes through better spatial–temporal profiling. We also plan to extend the benchmark to denser multi-entity settings and stronger camera motion. Finally, expanding the training data with more diverse positive and negative examples may further reduce hallucinations and improve coverage of more challenging physics understanding tasks.
Supplementary Material
Appendix A Additional Details of MASS-Bench
This section supplements the dataset description in Section 3 with expanded definitions and statistics.
A.1 Dataset Composition and Metadata
MASS-Bench aggregates three physics-centric video sources: 1,2291{,}229 samples ($14.7\%$) from MotionSight [17], 4,0094{,}009 ($47.9\%$) from VideoPhy2 [5], and 2,8622{,}862 ($34.2\%$) from VideoHallu [36]. In total, the dataset comprises 6,0936{,}093 training and 2,2682{,}268 test examples drawn from 4,3504{,}350 unique videos. Videos average $545.8$ frames (${\sim}19.6$ s at $27.4$ FPS) with a mean resolution of $1120{\times}702$ pixels. Due to overlap inherited from the source datasets, approximately $1.5\%$ of unique videos appear in both splits; however, no (video, question) pairs are shared across splits, so there is no direct QA leakage. All methods are evaluated on the same fixed test set with identical evaluation scripts. Importantly, MASS relies only on depth and tracking signals extracted from the input video at inference time and does not use any ground-truth test annotations.
A.2 Question Categories
We group all questions into five categories reflecting the type of physical reasoning required, ordered from basic perception to advanced comprehension. Table 5 shows the distribution.
Category Count Pct.
SU (Spatial Understanding) 2,785 33.3%
TU (Temporal Understanding) 1,633 19.5%
PA (Physical Abnormality Detection) 1,432 17.1%
PC (Physics Comprehension) 1,304 15.6%
MAR (Motion & Action Recognition) 1,205 14.4%
Table 5: Distribution of question categories in MASS-Bench.
(a) Spatial Understanding (SU). Identifying objects and their geometric relationships, positions, and scene layouts.
(b) Temporal Understanding (TU). Interpreting how events evolve over time, including ordering, duration, and temporal dependencies.
(c) Motion & Action Recognition (MAR). Detecting and characterizing object motions and agent actions across frames.
(d) Physics Comprehension (PC). Applying physical principles to infer, explain, or predict real-world dynamics.
(e) Physical Abnormality Detection (PA). Identifying motions or events that violate physical laws or exhibit implausible behavior.
These categories are ordered from easy to challenging. VLMs must first establish spatial and temporal awareness, then recognize motion patterns, before progressing to physics comprehension and violation detection.
A.3 Example and Question Types
We categorize each video–question–answer pair along two complementary axes.
For example type, we distinguish positive examples (3,4363{,}436, $41.1\%$), where the video follows real-world physics, from negative examples (4,9254{,}925, $58.9\%$), where at least one part of the scene violates physical plausibility. Negative examples are more challenging because the model must rely on visual evidence rather than defaulting to language priors.
For question type (Figure 2), we separate factual questions (5,4275{,}427, $67.0\%$) from critical-thinking questions (2,6732{,}673, $33.0\%$). Factual questions provide explicit cues that focus attention on particular entities or events. Critical-thinking questions are less direct: they omit such cues and instead require the model to infer intent, identify salient events, or reason about the underlying physical dynamics.
A.4 Motion Grounding Annotations
Beyond the question, video, and ground-truth answer, each pair includes five layers of motion-grounding annotation: (1) temporal video segmentation indexed by frame ranges; (2) visual grounding for each queried entity, specified by entity name and bounding box; (3) an entity-level temporal profile that tracks each grounded entity across the full video, persisting once detected and left blank in segments where the entity is absent; (4) per-segment motion attributes, including the first and last observed 3D positions to capture coarse spatial layout; and (5) 3D motion vectors representing each entity’s temporal displacement. Together, these cues convert physics-intensive perceptual challenges into structured textual and mathematical representations that support more reliable physical reasoning.
Task Description: The model receives a system instruction enforcing explicit reasoning and final answer formatting. Given a video, a question, and motion-grounding metadata, the model must produce detailed reasoning inside <think> tags and a concise final answer inside <answer> tags. Core Requirements: Use natural internal dialogue in the <think> section (e.g., “let me think”, “hmm”, “wait”). Perform step-by-step reasoning validating spatial–temporal cues. Place the final answer only inside <answer> tags. Use free-form answer format: Provide a short text answer within the answer tags.   Motion-Grounding Information: We use the following template to represent the motion-grounding information generated for VLMs. This template is used to fill the {motion_grouding_info} in the QA prompt template below: Entity #1: <Entity Name>
* Segment #1  First Position <first_position>, Motion Vector <motion>, Last Position <last_position>, Bounding Box <bbox>, Frame <first_frame>...<last_frame>
* Segment #2  First Position <first_position>, Motion Vector <motion>, Last Position <last_position>, Bounding Box <bbox>, Frame <first_frame>...<last_frame>
Entity #2: <Entity Name>
  Actual Prompt Used for Video QA: Conversation Setup:
A conversation between User and Assistant. The user asks a question, and the Assistant solves it. The Assistant first thinks about the reasoning process inside <think> </think> tags, then provides the final answer inside <answer> </answer> tags.
Question:
<Question> {question} </Question>
Motion-Grounding Information:
{motion_grouding_info}
Reasoning Instruction:
Please think about this question as if you were a human pondering deeply. Use internal dialogue such as ‘‘let me think’’, ‘‘wait’’, ‘‘hmm’’, ‘‘I see’’, and include verification or self-reflection in the reasoning process. Provide detailed reasoning in <think> </think>, then provide the final answer in <answer> </answer>.
Free-form Answer Instruction:
Please provide your text answer within the <answer> </answer> tags.
Prompt Template for Video-Language Models
Figure 4: Prompt template used for motion-aware video question answering. The template first serializes entity-level motion grounding (positions, motion vectors, bounding boxes, and frame ranges) into text, then injects this context into a chain-of-thought style prompt that guides the VLM to reason in <think> tags and output its final prediction in standardized <answer> tags.
Task Description: You are an intelligent teacher whose task is to evaluate the correctness of a model’s answer to a question, given a reference ground-truth answer. Inputs: Question: wrapped in <Question> ... </Question> Ground-truth answer: wrapped in <GT> ... </GT> Model prediction: wrapped in <Answer> ... </Answer> Evaluation Criteria: If the prediction does not conflict with the ground truth, output <Eval> Correct </Eval>. If the prediction conflicts with the ground truth, output <Eval> Incorrect </Eval>. If the correctness of the prediction is unclear, output <Eval> Unclear </Eval>. Reason carefully about the relationship between the prediction and the ground truth, but keep the final evaluation very brief. Output Format: Produce only one of the following tokens as the final output: <Eval> Correct </Eval> <Eval> Incorrect </Eval> <Eval> Unclear </Eval> Prompt Template for Evaluation
Figure 5: Prompt template used for automatic evaluation of model answers against ground-truth references. The template presents the question, ground truth, and model output provided for LLM-as-a-judge evaluation and guides the evaluator to produce one of three outcomes, Correct, Incorrect, or Unclear, ensuring reliable and consistent scoring across all predictions.
Appendix B Implementation Details
All experiments, including ablations, are conducted using full-parameter fine-tuning on 8 NVIDIA H100 (80GB) GPUs. Both Qwen2.5-VL-7B and LLaVA-OneVision-7B are trained under identical settings for fair comparison. The GRPO post-training phase for each model requires approximately 9–12 hours, while the supervised fine-tuning (SFT) stage in our ablation studies completes within 2–3 hours.
Appendix C Visual Feature Representation Template
Figure 4 illustrates the prompt template used during both post-training and inference for VLMs in MASS-Bench. This template is designed to elicit structured reasoning for free-form video question answering. The motion-grounding information incorporated into the prompt is derived from the spatial–motion feature extraction module and the visual representation pipeline described in Section 4.
For each detected entity, denoted as <Entity Name>, we generate motion-grounding descriptors for every video segment in which the entity appears. These descriptors include the entity’s first and last 3D positions within the segment (<first_position> and <last_position>), the corresponding 3D motion vector (<motion>), the bounding box in the segment’s first frame (<bbox>), and the segment’s temporal extent indicated by the starting and ending frame indices (<first_frame> and <last_frame>).
Appendix D Evaluation Template
We provide the LLM-as-a-judge evaluation template used in our experiments in Figure 5. This template standardizes how we assess the correctness of model predictions by comparing each VLM’s answer against the ground-truth annotation and the corresponding question.
Appendix E Experiments on Real-world Video QA
As a supplement to the experiments in the main paper, we provide additional results on two real-world video question-answering benchmarks, MMVU [67] and MVBench [30], to further assess the generalization ability of VLMs augmented with MASS. MMVU evaluates models on expert-level temporal, procedural, and interaction-centric reasoning, while MVBench is a widely used benchmark for temporal video understanding and action-centric tasks. Both datasets contain diverse video-based QA pairs captured from real-world scenarios.
Table 6 reports the performance of VLMs integrated with MASS alongside state-of-the-art baselines. On MMVU, large-scale close-source models achieve substantially higher accuracy due to the benchmark’s emphasis on expert-level, interdisciplinary reasoning—highlighting the limitations of smaller 7B models. Qwen2.5-VL-7B+MASS improves over its baseline on both benchmarks and narrows the gap to close-source models, while LLaVA-OneVision-7B+MASS shows a performance drop on these out-of-distribution tasks. As also shown in Table 8, LLaVA is more sensitive to detection and tracking quality: when salient entities are missed or motion cues are noisy—common in MMVU and MVBench—performance can degrade, motivating more selective motion-profile extraction in future work.
Model MMVU (%) MVBench (%)
Baselines
InternVL3.5-8B-Flash 55.04 45.78
Qwen2.5-VL-7B 64.00 51.30
LLaVA-OneVision-7B 50.41 55.24
Larger / Close-Source
GPT-4o 75.96 58.50
Gemini-2.5-Flash 75.27 54.32
VLMs with MASS
LLaVA-OneVision-7B + MASS 48.95 48.33
Qwen2.5-VL-7B + MASS 65.76 61.59
Table 6: Overall accuracies (%) on MMVU and MVBench. We report overall performance on two real-world video question-answering benchmarks, MMVU [67] and MVBench [30]. GPT-4o achieves the highest accuracy on both datasets. Qwen2.5-VL-7B+MASS improves over its baseline and narrows the gap to close-source models, while LLaVA-OneVision-7B+MASS is more sensitive to detection quality on these out-of-distribution tasks.
Appendix F Noise-Induced Robustness
We inject two noise types into the motion profile to test spatial–motion grounding robustness: (1) BBox/position noise, perturbing bounding boxes and initial/final 3D positions (SAM-2 & Grounding-DINO); (2) Trajectory Noise, perturbing 3D motion vectors (Co-Tracker). We apply a relatively large 10% Gaussian perturbation to the motion vectors, positions, and bounding boxes, both separately and jointly. Table 7 shows slight accuracy drops, yet MASS-augmented backbones still substantially outperform their baselines (Table 2 in the main paper) across all noise settings, even after RFT. Because our physical comprehension is mostly qualitative and the motion profiles serve as supplementary cues rather than exact measurements, the model tolerates moderate value-level perturbations. Failures mainly occur when noise flips the sign of motion components, i.e., reversing positive/negative values, leading the model to make wrong assumptions about motion directions and further cascading into a misinterpretation of the entire physical process.
Baselines (MASS) Noise Injection
Model (+ MASS) w/ w/o BBox Motion All
Qwen2.5-VL-7B 61.93 52.30 58.64 58.51 58.42
LLaVA-OneVision-7B 59.24 45.79 54.50 54.19 54.19
Table 7: Noise-induced robustness evaluation on MASS-augmented models. BBox, Motion, and All columns report accuracy under 10% Gaussian perturbations applied to bounding-box/position, 3D trajectory, and both components jointly. Baseline columns (w/ and w/o MASS) are provided as reference.
Appendix G Inference Time and Memory Usage
We report wall-clock inference time on 100 randomly sampled test cases on an NVIDIA L40S (Table 8). MASS increases per-query latency due to motion-profile extraction, while GPU memory increases only modestly. Our goal is to evaluate the benefit of explicit spatial–motion cues for physical understanding, not an optimized pipeline. Importantly, the motion profile is modular and cacheable per video, so it can be reused across multiple questions without re-extraction. Overhead can be further reduced via faster upstream modules and standard efficiency knobs (e.g., lower frame stride/resolution, smaller top-$K$ grounded entities), which we leave to future work.
Model Time (s/case) GPU Mem (GB)
w/o w/ w/o w/
LLaVA-OV-7B 1.86 6.49 14.97 15.69
Qwen2.5-VL-7B 5.83 10.46 15.45 16.18
Table 8: Per-query inference time (s) and GPU memory usage (GB) on an NVIDIA L40S over 100 randomly sampled test cases, with and without MASS.
Appendix H Multi-object Tracking and Camera Motion
We agree that dense tracking in crowded scenes increases compute and may introduce failures. MASS-Bench is entity-centric: we track only grounded target entities (top-$K$, tunable; default 5), and our QA design and video filtering avoid highly crowded, many-entity cases. Camera ego-motion remains an open challenge; the current benchmark largely contains limited camera motion to focus on entity-centric spatial/motion grounding. MASS does not assume perfect tracks and degrades gracefully under noisy motion cues (as shown in the noise robustness analysis in Table 7). We plan to extend the benchmark and method to crowded multi-entity tracking and stronger camera motion handling in future work.
Appendix I Case Study
In this section, we present additional qualitative examples of video question answering produced by several SoRA VLMs across the categories defined in MASS-Bench, as shown in Figures 610. Hallucinated or incorrect predictions are highlighted in red. Each example is accompanied by expert human annotations describing the ground-truth physics-driven dynamics, providing a clear reference for evaluating model behavior and identifying failure modes.
Figure 6: Video question-answering example from the Spatial Understanding (SU) category. We present physics reasoning and comprehension cases from state-of-the-art VLMs evaluated on spatial understanding tasks. Each example includes the video-generation prompt and human expert annotations, with visual grounding annotated (Red), the corresponding questions (Purple), and model responses from GPT-4o (Orange), Gemini-2.5-Flash (Blue), Qwen2.5-VL (Green), and Qwen2.5-VL + MASS (Gray). Hallucinated content and critical contextual errors are highlighted in Red.
Figure 7: Video question-answering example from the Temporal Understanding (TU) category. We present physics reasoning and comprehension cases from state-of-the-art VLMs evaluated on temporal understanding tasks. Each example includes the video-generation prompt and human expert annotations, with visual grounding annotated (Red), the corresponding questions (Purple), and model responses from GPT-4o (Orange), Gemini-2.5-Flash (Blue), Qwen2.5-VL (Green), and Qwen2.5-VL + MASS (Gray). Hallucinated content and critical contextual errors are highlighted in Red.
Figure 8: Video question-answering example from the Motion and Action Recognition (MAR) category. We present physics reasoning and comprehension cases from state-of-the-art VLMs evaluated on motion and action recognition tasks. Each example includes the video-generation prompt and human expert annotations, with visual grounding annotated (Red), the corresponding questions (Purple), and model responses from GPT-4o (Orange), Gemini-2.5-Flash (Blue), Qwen2.5-VL (Green), and Qwen2.5-VL + MASS (Gray). Hallucinated content and critical contextual errors are highlighted in Red.
Figure 9: Video question-answering example from the Physics Comprehension (PC) category. We present physics reasoning and comprehension cases from state-of-the-art VLMs evaluated on physics comprehension tasks. Each example includes the video-generation prompt and human expert annotations, with visual grounding annotated (Red), the corresponding questions (Purple), and model responses from GPT-4o (Orange), Gemini-2.5-Flash (Blue), Qwen2.5-VL (Green), and Qwen2.5-VL + MASS (Gray). Hallucinated content and critical contextual errors are highlighted in Red.
Figure 10: Video question-answering example from the Physics Abnormality Detection (PA) category. We present physics reasoning and comprehension cases from state-of-the-art VLMs evaluated on physics abnormality detection tasks. Each example includes the video-generation prompt and human expert annotations, with visual grounding annotated (Red), the corresponding questions (Purple), and model responses from GPT-4o (Orange), Gemini-2.5-Flash (Blue), Qwen2.5-VL (Green), and Qwen2.5-VL + MASS (Gray). Hallucinated content and critical contextual errors are highlighted in Red.